### v1

@@ Abstract
The stability of plasmas is a challenging topic. We study a control problem for plasma living in a uniform external magnetic field. Linear analysis for any equilibrium $\mu$ is investigated through the Laplace-Fourier transform approach. Moreover, the Penrose condition, which characterizes the stability of equilibrium, is derived for Bernstein modes and verified numerically. Based on the linear analysis, a general control strategy is introduced, with recovering the free-streaming solution as a specific example. We finally present several examples to control instabilities, including Gaussian equilibria and Dory-Guest-Harris instability.
@@
Plasma, Vlasov-Poisson equation, Control, Bernstein modes, Dory-Guest-Harris instability, External electric field
@@ 1 Introduction
Plasma is known as the fourth state of matter, which is an ionized gas consisting of free electrons and ions, and is also the main constituent of ordinary matter. It holds the promise of fusion and is important to understand many astrophysical phenomena. However, plasma can be highly unstable even when slightly perturbed from equilibrium, which poses challenges to experiments and applications [1, 2]. This is due to complex interactions between particles and fields at the microscopic level. Therefore, controlling plasma has attracted increasing attention.
We study a plasma placed in an external magnetic field. One such example is the LAPD [3] device with straight magnetic field lines. Additional examples include magnetic confinement fusion reactors such as tokamaks, stellarators, and mirrors. In this scenario, the plasma dynamics can be modeled by the Vlasov-Poisson (VP) system with a uniform external magnetic field. Denoting the electron distribution $F(t,\mathbf{x},\mathbf{v})$, the self-generated electric potential $V(t,\mathbf{x})$ and the associated self-generated electric field $\mathbf{E}(t,\mathbf{x})$, VP system writes:
@@
where $\mathbf{x}=(x_{1},x_{2},x_{3})^{T}\in\mathbb{T}^{3},\mathbf{v}=(v_{1},v_{2},v_{3})^{T}\in\mathbb{R}^{3}$ are the spatial and velocity coordinates. $\mathbf{B}=(0,0,B_{0})^{T}$ is the fixed external magnetic field pointing in the $x_{3}$-direction, with $B_{0}>0$. For notational convenience, throughout the paper, we denote $\mathbf{x}_{\perp}:=(x_{1},x_{2})^{T},\mathbf{v}_{\perp}:=(v_{1},v_{2})^{T}$.
The VP system with an external magnetic field demonstrates features that were unseen to those without an external magnetic field in previous studies [4, 5, 6]. The foremost discovery is that the structure of equilibrium distributions has changed. Due to the impact of the external magnetic field, only distribution that has radial symmetry in $\mathbf{v}_{\perp}=(v_{1},v_{2})$ space can serve as equilibrium states. Throughout the paper, we write our equilibrium state as a separable function $\mu(\mathbf{v})=\mu_{\perp}(\mathbf{v}_{\perp})\mu_{\parallel}(v_{3})$. As in the classical VP stability analysis, some equilibrium states are stable while others are not, depending on the profile of $\mu_{\perp}$ and $\mu_{\parallel}$. In this paper, we place focus on two kinds of equilibrium states with $\mu_{\perp}$ being either a Gaussian or ring-shaped. The former was shown in [7] to have non-damping components of electric energy, and the latter can form Dory-Guest-Harris instability [8, 9, 10]. Another feature of the VP system with external magnetic field is the concept of Bernstein modes, the collection of modes corresponding to $\mathbf{k}=(k_{1},k_{2},0)^{T}$. Since $\mathbf{B}$ has its field line pointing in $x_{3}$-direction, it is of particular importance to investigate the stability of the solution on the $\mathbf{x}_{\perp}=(x_{1},x_{2})$ plane that is perpendicular to the field line, and such stability is uniquely characterized by Bernstein modes. As a consequence, we confine ourselves to these modes in numerical simulations.
@@
In magnetic confinement fusion, modes with $k_{3}=0$ are routinely investigated in order to assess the relative importance of drift-waves, the Kelvin-Helmholtz (KH) instability, and sheath physics in driving turbulence and setting its saturation levels [3]. Experimentally [11, 12], plasma transport in the Large Plasma Device (LAPD) due to KH modes with $k_{3}=0$ is studied using an external biasing of the plasma relative to the chamber wall. Additionally, in the Mirabelle experiment [13], strong KH mode activity is obtained for various source profiles. Theoretically, the approximation $k_{3}\ll k_{1},k_{2}$ is typically employed in magnetized plasma systems due to the fact that the strong guide field leads to a separation of scales, with turbulence in $x_{1}$ and $x_{2}$ occurring on the gyroradius scale and structures along $x_{3}$ being on the order of the macroscopic length $L$. These are the approximations used to model magnetized plasmas using either a fluid framework (e.g., MHD and drift-reduced Braginskii equations) or a kinetic one (e.g., drift-kinetic and gyrokinetic equations). Finally, besides KH and sheath modes, we mention the Gradient-driven Drift Coupling (GDC) instability, which, as an essentially $k_{3}\simeq 0$ mode, is regarded as a candidate mechanism for turbulence in basic, space, and astrophysical plasmas [14].
@@
From the mathematical side, the Landau damping result was conducted in [15] where the authors developed the general recipe for studying the PDE spectrum for the VP system. When an external magnetic field is present, in [7], the authors derived the dispersion relation for the linearized VP system and investigated the Landau damping for Bernstein modes when $\mu_{\parallel}$ is a Gaussian. Extending from their work, we removed the Gaussian assumption on $\mu$ and achieved a more general dispersion relation.
The control problem is also not new. Such a control problem is more prominently studied in the context of Magnetohydrodynamic (MHD) system [2]. In the kinetic framework, a numerical investigation was conducted for the classical VP system in $1D1V$ setting in [16] using an external electric field. This work is complemented by [17, 18, 19, 20], where the authors achieve control using an external magnetic field. These numerical studies were then followed by theoretical derivation in [21], where the authors proposed the pole elimination strategy as a general method for suppressing instability. Following these footsteps, we adopt the pole elimination principle and apply it to (1.1) for the design of $H$. This design essentially eliminates the positive eigenvalues in the PDE, and thus removes the growing components from the system, achieving the suppression of instability.
@@ 2 Linear analysis
$\partial_{t}f+\mathbf{v}\cdot\nabla_{\mathbf{x}}f+(v_{2}B_{0}\partial_{v_{1}}f-v_{1}B_{0}\partial_{v_{2}}f)+\mathbf{E}\cdot\nabla_{\mathbf{v}}\mu(\mathbf{v})$ $=0\,$, (2.2)
@@
We first perform the derivation of the dispersion relation for (2.2). The technique has been deployed in [15] for Vlasov-Poisson, a slightly simpler system.
@@
For this linear ODE system, the solution is explicit:
@@
we have, by plugging in (2.4):
@@
$=\hat{g}(t,\mathbf{k}_{\perp},k_{3},(\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp},k_{3}t)\,$, (2.19)
@@
$\dfrac{\mathrm{d}\hat{g}}{\mathrm{d}t}(t,\mathbf{k},\mathbf{m})=$ $-\int_{\mathbb{T}^{3}}\int_{\mathbb{R}^{3}}e^{-i\mathbf{k}\cdot\mathbf{x}-i\mathbf{m}\cdot\mathbf{v}}\mathbf{E}(\mathbf{y}(t;\mathbf{x},\mathbf{v}))\cdot\nabla_{\mathbf{v}}\mu(\mathbf{u}(t;\mathbf{v}))\,d\mathbf{x}d\mathbf{v}$ (2.20)
@@
$=$ $-i\begin{pmatrix}\hat{\mathbf{E}}_{\perp}(t,\mathbf{k})\\ \hat{E}_{3}(t,\mathbf{k})\end{pmatrix}\cdot\begin{pmatrix}\mathbf{R}(B_{0}t)(\mathbf{m}_{\perp}-(\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp})\\ m_{3}-k_{3}t\end{pmatrix}$
$\hskip 79.49744pt\hat{\mu}(\mathbf{R}(B_{0}t)(\mathbf{m}_{\perp}-(\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp}),m_{3}-k_{3}t)\,.$
This equation holds for all values of $(\mathbf{k},\mathbf{m})$. We chose a special pair. Let $\mathbf{m}=(\mathbf{m}_{\perp},m_{3})$ defined by:
$\mathbf{m}_{\perp}(\mathbf{k}):=(\mathbf{M}(t)+\mathbf{Q})^{T}\mathbf{k}_{\perp}\,,\quad m_{3}(\mathbf{k}):=k_{3}t\,$, (2.21)
@@
$\mathbf{R}(B_{0}s)(\mathbf{m}_{\perp}(\mathbf{k})-(\mathbf{M}(s)+\mathbf{Q})^{T}\mathbf{k}_{\perp})$ $=\mathbf{R}(B_{0}s)(\mathbf{M}(t)-\mathbf{M}(s))^{T}\mathbf{k}_{\perp}\,$,
@@
hence proving (2.5). The stability condition (2.9) comes from the direct application of the Mellin inversion theorem for inverting the Laplace transform. ∎
@@ 2.2 Two case studies
In this subsection, we compute (2.6) for two equilibria, the Gaussian equilibrium and a ring-shaped equilibrium that generates Dory-Guest-Harris instability. Moreover, we restrict ourselves to the Bernstein modes, the Fourier modes of $\mathbf{k}=(k_{1},k_{2},0)^{T}$ that are transverse to the magnetic field. Consequently, the derivation of the rest of the paper is conducted in $2D2V$.
@@
For another specific mode in the form of $(0,0,k_{3})^{T}$, our linear analysis recovers the previous results of Vlasov-Poisson [21], where $\mu_{\parallel}(v_{3})$ is two-stream instability or bump-on-tail instability along the magnetic field direction.
@@
We first consider Gaussian equilibrium of (2.2). The Gaussian equilibrium has been proved to have non-decaying electric energy for Bernstein modes [7]. That is, for each $\mathbf{k}=(k_{1},k_{2})^{T}\in\mathbb{Z}^{2}$, the poles of $P(\mathbf{k},\lambda)$ are located along the imaginary line.
Let $\mu$ be any isotropic Gaussian distribution on $(v_{1},v_{2})^{T}$, i.e.
@@
$\mathbf{k}\cdot\tilde{\mathbf{R}}(t)\mathbf{k}$ $=\begin{pmatrix}k_{1}\\ k_{2}\end{pmatrix}^{T}\frac{1}{B_{0}}\begin{pmatrix}\sin(B_{0}t)&\cos(B_{0}t)-1\\ -\cos(B_{0}t)+1&\sin(B_{0}t)\end{pmatrix}\begin{pmatrix}k_{1}\\ k_{2}\end{pmatrix}=\frac{\sin(B_{0}t)}{B_{0}}|\mathbf{k}|^{2}\,.$
In the following we set $\theta=1$, $B_{0}=1$ and $\mathbf{k}_{\perp}=(1,0)^{T}$ without loss of generality, then
@@
In Fig. 1, we plot the equilibrium (2.27) and the evaluation of $P(\mathbf{k},\lambda)$ for $\lambda=\mathrm{Re}(\lambda)+i\mathrm{Im}(\lambda)$ with $(\mathrm{Re}(\lambda),\mathrm{Im}(\lambda))\in[0.001,0.1]\times[1.05,1.25]$. Since poles for $P$ are complex conjugate pairs, we only plot them on the $\mathrm{Im}(\lambda)>0$. The plot shows that we can numerically find a root at $(\mathrm{Re}\lambda=0.001,\mathrm{Im}(\lambda)=1.164)$. This is consistent with the analytical prediction in [7] that suggests all roots should be along the imaginary line [^1] Numerically, we cannot compute $P(\mathbf{k},\lambda)$ for $\mathrm{Re}(\lambda)=0$..

### v2

@@ Abstract
Stabilizing plasma dynamics through externally applied electric and magnetic fields is a fundamental control problem. We study this question for a plasma evolving under a uniform external magnetic field. Although the governing dynamics are nonlinear, a linear analysis based on the Laplace–Fourier transform yields actionable insight. In particular, by controlling the location of the roots of the dispersion relation, we propose a general control strategy that restores stability, with the free-streaming solution recovered as a special case. Numerical experiments for Gaussian equilibria and for the Dory–Guest–Harris instability show that the proposed control suppresses the unstable modes and stabilizes the dynamics, in agreement with our theoretical predictions.
@@
Plasma, Vlasov-Poisson equation, Control, Bernstein modes, Dory-Guest-Harris instability, External electric field, Uniform external magnetic field
@@ 1 Introduction
Plasma is often referred to as the fourth state of matter: an ionized gas consisting of free electrons and ions. It is ubiquitous in nature and plays a central role in astrophysical phenomena, and it also underlies the promise of controlled nuclear fusion. However, plasma dynamics can be highly unstable. Near certain equilibria, even small perturbations may trigger rapid growth of instabilities and drive the system into strongly nonlinear, possibly turbulent, regimes, posing major challenges for both experiments and applications [Solovev-Shafranov-1995-mag-confinement, zohm2015magnetohydrodynamic, Freidberg_2007]. In fusion devices, for instance, one aims to maintain the plasma close to a stable equilibrium for as long as possible so that the energy remains primarily in kinetic form. Instabilities can instead induce a fast transfer of energy into electric fields and other channels, making energy extraction difficult. Motivated by these considerations, we study plasma control strategies aimed at stabilizing the dynamics near equilibrium.
In most experimental configurations, such as tokamaks [Wesson_tokamak_2004], stellarators, and mirror machines, plasma is confined by an externally applied magnetic field. We focus on a regime in which the magnetic field lines are approximately straight, a setting exemplified by the LAPD (LArge Plasma Device) at UCLA [Rogers-Ricci-LAPD-2010]. In this case, the plasma dynamics can be modeled by the Vlasov–Poisson (VP) system under a uniform external magnetic field. Denoting the electron distribution $F(t,\mathbf{x},\mathbf{v})$, the self-generated electric potential $V(t,\mathbf{x})$ and the associated self-generated electric field $\mathbf{E}(t,\mathbf{x})$, VP system writes:
@@
where $\frac{dF}{dt}$ denotes the total time variation of $F$, $\mathbf{x}=(x_{1},x_{2},x_{3})^{\top}\in\mathbb{T}^{3},\mathbf{v}=(v_{1},v_{2},v_{3})^{\top}\in\mathbb{R}^{3}$ are the spatial and velocity coordinates. $\mathbf{B}=(0,0,B_{0})^{\top}$ is the fixed external magnetic field pointing in the $x_{3}$-direction, with $B_{0}>0$. For notational convenience, throughout the paper, we denote $\mathbf{x}_{\perp}:=(x_{1},x_{2})^{\top},\mathbf{v}_{\perp}:=(v_{1},v_{2})^{\top}$.
The added external magnetic field brings features to the VP system compared to the one without them [Schaeffer-1991-vlasov-poisson, PFAFFELMOSER1992281, Mouhot-Villani-2011-landau, Skubachevskii2015-zl, Belyaeva_2020_VP_cylinder, Skubachevskii_2021_VP_confine] . The foremost discovery is that the structure of equilibrium distributions has changed. Due to the impact of the external magnetic field, only distribution that has radial symmetry in $\mathbf{v}_{\perp}=(v_{1},v_{2})$ space can serve as equilibrium states. Throughout the paper, we write our equilibrium state as a separable function $\mu(\mathbf{v})=\mu_{\perp}(\mathbf{v}_{\perp})\mu_{\parallel}(v_{3})$. As in the classical VP stability analysis, some equilibrium states are stable while others are not, depending on the profile of $\mu_{\perp}$ and $\mu_{\parallel}$. In this paper, we place focus on two kinds of equilibrium states with $\mu_{\perp}$ being either a Gaussian or ring-shaped. The former was shown in [Bedrossian-Wang-2020-VP] to have non-damping components of electric energy, and the latter can form Dory-Guest-Harris instability [DGH-1965-PRL, Tataronis-Crawford-1970-perpendicular, VOGMAN2014101]. Another characteristic feature of the Vlasov–Poisson system in the presence of a uniform magnetic field is the appearance of Bernstein modes [Bernstein-1958, Sukhorukov-1997-Bernstein, Bedrossian-Wang-2020-VP]. When $\mathbf{B}$ is directed along the $x_{3}$-axis, it is natural to distinguish the dynamics parallel and perpendicular to the field. Bernstein modes correspond to electrostatic waves with perpendicular wavevector, i.e. $k_{\parallel}=0$; in our notation this means $\mathbf{k}=(k_{1},k_{2},0)^{\top}$. This restriction places the focus on stability properties in the transverse plane $\mathbf{x}_{\perp}=(x_{1},x_{2})$, where magnetization induces behavior distinct from the unmagnetized case. As a consequence, we confine ourselves to these modes in numerical simulations.
@@
To understand plasma dynamics, modes with $k_{3}=0$ are routinely investigated in order to assess the relative importance of drift-waves, the Kelvin-Helmholtz (KH) instability, and sheath physics in driving turbulence and setting its saturation levels [Rogers-Ricci-LAPD-2010]. Experimentally [Horton-etal-PhysPlasma-2005-KH, Perez-etal-PhysPlasma-2006-LAPD], plasma transport in LAPD due to KH modes with $k_{3}=0$ is studied using an external biasing of the plasma relative to the chamber wall. Additionally, in the Mirabelle experiment [Brochard-etal-PhysPlasma-2005-KH], strong KH mode activity is obtained for various source profiles. Theoretically, the approximation $k_{3}\ll k_{1},k_{2}$ is typically employed in magnetized plasma systems due to the fact that the strong guide field leads to a separation of scales, with turbulence in $x_{1}$ and $x_{2}$ occurring on the gyroradius scale and structures along $x_{3}$ being on the order of the macroscopic length $L$. These are the approximations used to model magnetized plasmas using either a fluid framework (e.g., MHD and drift-reduced Braginskii equations) or a kinetic one (e.g., drift-kinetic and gyrokinetic equations). Finally, besides KH and sheath modes, we mention the Gradient-driven Drift Coupling (GDC) instability, which, as an essentially $k_{3}\simeq 0$ mode, is regarded as a candidate mechanism for turbulence in basic, space, and astrophysical plasmas [Pueschel-etal-PlasmaPhys-2017-LAPD].
@@
A cornerstone to understand plasma dynamics is to derive and analyze the dispersion relation. This is closely tied to the PDE theory on the study of the operators’ spectrum. In the context of VP system, a celebrated work on Landau damping was rigorously proved in [Bedrossian-Masmoudi-Mouhot-2016-Landau]. The dispersion relation for the system with an external magnetic field was studied in [Bedrossian-Wang-2020-VP], where the authors investigated the stability of the Bernstein modes for $\mu_{\parallel}$ being a Gaussian. Extending from their work, we removed the Gaussian assumption on $\mu$ and achieved a more general dispersion relation.
The plasma control problem has also been extensively studied in the literature. It is particularly prominent in the magnetohydrodynamic (MHD) setting; see, e.g., [zohm2015magnetohydrodynamic]. In the kinetic framework, a numerical investigation of the classical VP system in the 1D1V setting using an externally applied electric field was carried out in [EINKEMMER2024112662]. Related numerical studies include [Knopf2018-fc, Knopf-Weber-vp-mag-control-2020-bw, Patrik_Knopf_control_PIC, ALBI2025113804], where control is achieved via externally applied magnetic fields. These numerical works were followed by the theoretical development in [EINKEMMER-Li-Mouhot-Yue-2025], which introduced the pole-elimination strategy as a general mechanism for suppressing instability. Building on this idea, we develop a pole-elimination–based control design for the magnetized Vlasov–Poisson system (1.1). Our construction uses a detailed analysis of the associated dispersion relation to derive an explicit feedback term $H$ that removes the unstable spectral contributions (i.e., modes with positive growth rates) and thereby stabilizes the dynamics. While conceptually inspired by [EINKEMMER-Li-Mouhot-Yue-2025], the present setting introduces substantial new technical challenges: the analysis is carried out in the full 3D3V phase space and must account for the anisotropic geometry induced by a strong external magnetic field, including the distinguished role of the $x_{3}$-direction. This leads to new phenomena absent in the unmagnetized case, such as the treatment of ring-type equilibria and the presence of Bernstein modes ($k_{3}=0$), both of which require additional structural considerations in the control design and stability analysis. We also note that this direction differs from work on magnetic confinement [Skubachevskii2015-zl, Belyaeva_2020_VP_cylinder, Skubachevskii_2021_VP_confine], which focuses on maintaining spatial confinement, instead of stability, of the plasma in bounded geometries under external fields.
@@ 2 Linear analysis
$\partial_{t}f+\mathbf{v}\cdot\nabla_{\mathbf{x}}f+(v_{2}B_{0}\partial_{v_{1}}f-v_{1}B_{0}\partial_{v_{2}}f)+\mathbf{E}\cdot\nabla_{\mathbf{v}}\mu(\mathbf{v})=0\,,\quad{\mathbf{x}\in\mathbb{T}^{3},\mathbf{v}\in\mathbb{R}^{3}}$ (2.2)
@@
We first perform the derivation of the dispersion relation for (2.2). The technique has been deployed in [Bedrossian-Masmoudi-Mouhot-2016-Landau] for Vlasov-Poisson, a slightly simpler system.
@@
Complementary to the Penrose condition (2.9), suppose that there exist $\mathbf{k}\in\mathbb{Z}^{3}$ and $\lambda_{0}\in\mathbb{C}$ such that $P(\mathbf{k},\lambda_{0})=0$, then the temporal behavior of the corresponding mode is characterized by the complex root $\lambda_{0}$. In particular, the real part of $\lambda_{0}$, $\mathrm{Re}(\lambda_{0})$, determines whether the mode grows, decays, or remains neutral in time. If $\mathrm{Re}(\lambda_{0})=0$, the associated electric energy does not grow or decay and corresponds to a neutrally stable mode. If $\mathrm{Re}(\lambda_{0})>0$, the mode exhibits exponential growth in time, indicating a potential instability.
@@
For this linear ODE system, the solution is explicit — the celebrated the Larmour trajectory (rotation) :
@@
we have, by plugging in the linearized equation of $f$ (2.2):
@@
$=\hat{g}(t,\mathbf{k}_{\perp},k_{3},(\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp},k_{3}t)\,$, (2.19)
@@
$\dfrac{\mathrm{d}\hat{g}}{\mathrm{d}t}(t,\mathbf{k},\mathbf{m})=$ $-\int_{\mathbb{T}^{3}}\int_{\mathbb{R}^{3}}e^{-i\mathbf{k}\cdot\mathbf{x}-i\mathbf{m}\cdot\mathbf{v}}\mathbf{E}(\mathbf{y}(t;\mathbf{x},\mathbf{v}))\cdot\nabla_{\mathbf{v}}\mu(\mathbf{u}(t;\mathbf{v}))\,d\mathbf{x}d\mathbf{v}\,$, (2.20)
@@
$=$ $-i\begin{pmatrix}\hat{\mathbf{E}}_{\perp}(t,\mathbf{k})\\ \hat{E}_{3}(t,\mathbf{k})\end{pmatrix}\cdot\begin{pmatrix}\mathbf{R}(B_{0}t)(\mathbf{m}_{\perp}-(\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp})\\ m_{3}-k_{3}t\end{pmatrix}$
$\hskip 79.49744pt\hat{\mu}(\mathbf{R}(B_{0}t)(\mathbf{m}_{\perp}-(\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp}),m_{3}-k_{3}t)\,.$
This equation holds for all values of $(\mathbf{k},\mathbf{m})$. We choose a special pair. Let $\mathbf{m}=(\mathbf{m}_{\perp},m_{3})$ defined by:
$\mathbf{m}_{\perp}(\mathbf{k}):=(\mathbf{M}(t)+\mathbf{Q})^{\top}\mathbf{k}_{\perp}\,,\quad m_{3}(\mathbf{k}):=k_{3}t\,$, (2.21)
@@
$\mathbf{R}(B_{0}s)(\mathbf{m}_{\perp}(\mathbf{k})-(\mathbf{M}(s)+\mathbf{Q})^{\top}\mathbf{k}_{\perp})$ $=\mathbf{R}(B_{0}s)(\mathbf{M}(t)-\mathbf{M}(s))^{\top}\mathbf{k}_{\perp}$
@@
hence proving (2.5). The stability condition (2.9) comes from inverting the Laplace transform [Davies_1978_transform]. ∎
@@ 2.2 Two case studies
In this subsection, we compute (2.6) for two equilibria, the Gaussian equilibrium and a ring-shaped equilibrium that generates Dory-Guest-Harris instability. Moreover, we restrict ourselves to the Bernstein modes, the Fourier modes of $\mathbf{k}=(k_{1},k_{2},0)^{\top}$ that are transverse to the magnetic field. Consequently, the derivation of the rest of the paper is conducted in $2D2V$.
@@
For another specific mode in the form of $(0,0,k_{3})^{\top}$, our linear analysis recovers the previous results of Vlasov-Poisson [EINKEMMER-Li-Mouhot-Yue-2025], where $\mu_{\parallel}(v_{3})$ is two-stream instability or bump-on-tail instability along the magnetic field direction.
@@
We first consider Gaussian equilibrium of (2.2). The Gaussian equilibrium has been proved to have non-decaying electric energy for Bernstein modes [Bedrossian-Wang-2020-VP]. That is, for each $\mathbf{k}=(k_{1},k_{2})^{\top}\in\mathbb{Z}^{2}$, the poles of $P(\mathbf{k},\lambda)$ are located along the imaginary line.
Let $\mu$ be any isotropic Gaussian distribution on $(v_{1},v_{2})^{\top}$, i.e.
@@
$\mathbf{k}\cdot\tilde{\mathbf{R}}(t)\mathbf{k}$ $=\begin{pmatrix}k_{1}\\ k_{2}\end{pmatrix}^{\top}\frac{1}{B_{0}}\begin{pmatrix}\sin(B_{0}t)&\cos(B_{0}t)-1\\ -\cos(B_{0}t)+1&\sin(B_{0}t)\end{pmatrix}\begin{pmatrix}k_{1}\\ k_{2}\end{pmatrix}=\frac{\sin(B_{0}t)}{B_{0}}|\mathbf{k}|^{2}\,.$
In the following we set $\theta=1$, $B_{0}=1$ and $\mathbf{k}_{\perp}=(1,0)^{\top}$ without loss of generality, then
@@
In Figure 1, we plot the equilibrium (2.27) and the evaluation of $P(\mathbf{k},\lambda)$ for $\lambda=\mathrm{Re}(\lambda)+i\mathrm{Im}(\lambda)$ with $(\mathrm{Re}(\lambda),\mathrm{Im}(\lambda))\in[0.001,0.1]\times[1.05,1.25]$. Since poles for $P$ are complex conjugate pairs, we only plot them on the $\mathrm{Im}(\lambda)>0$. The plot shows that we can numerically find a root at $(\mathrm{Re}\lambda=0.001,\mathrm{Im}(\lambda)=1.164)$. This is consistent with the analytical prediction in [Bedrossian-Wang-2020-VP] that suggests all roots should be along the imaginary line [^1] Numerically, we cannot compute $P(\mathbf{k},\lambda)$ for $\mathrm{Re}(\lambda)=0$..